Ace Industrial Machines issued 125,000 zero coupon bonds 9 years ago. The bonds originally had 30 years to maturity with a yield to maturity of 6 percent. Interest rates have recently decreased, and the bonds now have a yield to maturity of 5.1 percent. The bonds have a par value of $2,000. If the company has a $77.8 million market value of equity, what weight should it use for debt when calculating the cost of capital? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., .1616.)
Answer #1
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =21
Bond Price =∑ [(0*2000/100)/(1 + 6/100)^k]     +   2000/(1 + 6/100)^21
                   k=1
Bond Price = 588.31

MV of debt = price*bonds = 588.31*125000

=

73538750
Total Capital value = Value of Debt + Value of Equity
=73538750+77800000
=151338750
Weight of Debt = Value of Debt/Total Capital Value
= 73538750/151338750
=0.4859